OpenStudy (anonymous):

2+2*8-6x=19 \[x=-\frac{ 1 }{ 6 }\]

OpenStudy (secret-ninja):

You are solving for x, correct?

OpenStudy (haleyelizabeth2017):

PEMDAS

OpenStudy (haleyelizabeth2017):

And just to warn you, it does not come out to an even number, if I did it correctly.

OpenStudy (secret-ninja):

To solve for x, you have to get it by itself on the other side of the equal sign. Start by moving the "-6x" to the right. It will look like this: 2+2*8=19+6x You change it to +6x because the symbol becomes its opposite.

OpenStudy (haleyelizabeth2017):

According to the order of operations, you would need to multiply 2 by 8 first, then add and subtract.

OpenStudy (secret-ninja):

Now you move the 19 to the other side, and since it doesn't not have a symbol, it is automatically positive, and you change it to negative when moving it left. 2+2*8-19=6x Now solve for this: 2+2*8-19 2+16-19 18-19 -1 So you now have: -1=6x To get to the x, you divide because since the x is touching the 6 they are multiplying, and you are to do the opposite. Solve: 6 divided by -1 = x

OpenStudy (secret-ninja):

\[6\div-1=x\]

OpenStudy (secret-ninja):

What do you think x is?

OpenStudy (secret-ninja):

6 divided by -1 is -6, so x=-6

OpenStudy (secret-ninja):

\(\bbox[8pt, pink, border:5pt solid green]{\Large Welcome~to~Openstudy!~:)}\) If this helped, please click the \(\large\sf\color{blue}{Best~Response}\) button and then click the \(\large\sf\color{blue}{Close}\) button to close your question so that you can post a new one later. Go here for all of the basics: http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8 Go Here to learn about the rules of Os: http://openstudy.com/code-of-conduct \(\Huge\cal\color{darkgreen}{Have~a~great~day!}\)

OpenStudy (secret-ninja):

@chrisjones32

OpenStudy (haleyelizabeth2017):

it would actually be \[x=-\frac{ 1 }{ 6 }\] because 6x was equal to -1. then you would divide -1 by 6.